amm:

The points (3, 24) and (7, 56) represent points of a function where y, the number of photographs, varies directly with x, the number of pages in an album. Which statement describes another point on the graph of this function?

Maple:

so if we plug in the two points we have into the above equation we can solve of a and b. \(24 = a*3 + b // 56 = a*7 + b\) if \(24 = a*3 + b\), that means that \(b = 24 - a*3\), we can now substitute b in the other equation to get: \(56 = a*7 + 24 - a*3\) simplify this down we get: \(56 - 24 = a*4\) \(32 = a*4\) \(a = 8;\) put back a into any of the previous equations and we can solve for b \(24 = 8*3 + b\) \(b = 0;\) Now we have our equation/relation: \(y = 8x;\) we could have also looked at the points and realized that y is 8 times x.

amm:

thank you

Maple:

My pleasure, that's what I'm here for!